A student has test grades of 91,86 and 85,what must be her next test grade in order to have a 90 average
Found 2 solutions by Alan3354, greenestamps:
Answer by Alan3354(69443)

A student has test grades of 91,86 and 85,what must be her next test grade in order to have a 90 average
----------
G = the next grade
(91 + 86 + 85 + G)/4 = 90
Solve for G

Answer by greenestamps(13198)



The first grade, 91, is +1 compared to the desired average.
The second grade, 86, is -4 compared to the desired average.
The third grade, 85, is -5 compared to the desired average.

All together, the three grades are +1-4-5 = -8 compared to the desired average.

That means the next score has to be +8 compared to the desired average.

ANSWER: 90+8 = 98